OpenStudy (dtan5457):

Solve for x. x^5+8x^3-9x=0

OpenStudy (jhannybean):

Factor out the lowest common factor within all three terms first. Between \(x^5 ~, ~ x^3~,~ x^1\) which one is the smallest?

OpenStudy (dtan5457):

x^1

OpenStudy (jhannybean):

Alright, so factor out an \(x\) from the equation.

OpenStudy (dtan5457):

x(x^4+8x^2-9)

OpenStudy (dtan5457):

what happened to the 8x^2?

OpenStudy (jhannybean):

You are right, \(x^4 +8x^2 -9\).

OpenStudy (jhannybean):

Solve it like you would any regular quadratic. Treat your \(x^4\) like an \(x^2\)

OpenStudy (jhannybean):

So find two numbers that multiply to give -9, add to give +8. What are they?

OpenStudy (dtan5457):

-1, 9

OpenStudy (dtan5457):

im not too sure what you said about x^4+8x^2-9, how do i make this a regular quadratic?

OpenStudy (zzr0ck3r):

let \(y=x^2\) now we have \(y^2+8y-9=0\)

OpenStudy (zzr0ck3r):

can you solve this for y?

OpenStudy (dtan5457):

9 -1

OpenStudy (jhannybean):

Good, so if we use @zzr0ck3r 's method, letting \(y=x^2\), then we ca turn our quadfratic into it's factored form. \(y^2+8y-9=0 \iff (y+9)(y-1)=0\)

OpenStudy (jhannybean):

quadratic*

OpenStudy (jhannybean):

Now all we have to do is resubstitute \(x^2\) in for \(y\), and that'll leave us with: \[(x^2+9)(x^2-1)=0\]

OpenStudy (dtan5457):

is that -3 and i?

OpenStudy (zzr0ck3r):

\(\pm 3i\)

OpenStudy (jhannybean):

\[x(x^2+9)(x^2-1)=0\]\[x=0\]\[x^2=-9\implies x=\pm 3i\]\[x^2=1\implies x=\pm 1\]

OpenStudy (jhannybean):

More specifically, \(x^2 = \sqrt{-9} = \sqrt{-1} \cdot \sqrt{9} = i\cdot(\pm 3) = \pm 3i \)

OpenStudy (jhannybean):

always remember: \(i= \sqrt{-1}~,~ i^2 = -1\)

OpenStudy (dtan5457):

Very nicely explained. Thank you!

OpenStudy (jhannybean):

No problem :)